# The Coverage Principle: A Framework for Understanding Compositional Generalization

### v1

@@ D.3 Token set size ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig. 9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig. 10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
@@ D.4 Task ablation
Appendix E Derivation of Result 5.1 (power–law lower bound)
Problem setting
Let ${\mathcal{X}}$ be a finite token set with cardinality $|{\mathcal{X}}|$. The target mapping
$f:{\mathcal{X}}^{3}\longrightarrow{\mathcal{X}},\qquad f(x_{1},x_{2},x_{3})=f_{2}\bigl{(}f_{1}(x_{1},x_{2}),\,x_{3}\bigr{)}$
is a two-hop composition of unknown primitives $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. Write $b=f_{1}(x_{1},x_{2})\in{\mathcal{X}}$ for the intermediate state. Throughout the proof we impose two assumptions.
  1. (A1)
    Balanced classes: Each intermediate value $b\in{\mathcal{X}}$ is realized by exactly $|{\mathcal{X}}|$ first-hop pairs, i.e., the sets $E_{b}:=\{(x_{1},x_{2})\in{\mathcal{X}}^{2}:f_{1}(x_{1},x_{2})=b\}$ all have size $|{\mathcal{X}}|$ and form a partition of ${\mathcal{X}}^{2}$.
  2. (A2)
    Uniform training sampler: The training set $D$ contains $N$ triples drawn uniformly with replacement from the domain ${\mathcal{X}}^{3}$.
Functional $k$-equivalence and learner model
For two first-hop fragments $a,a^{\prime}\in{\mathcal{X}}^{2}$ define
$a\sim a^{\prime}\quad:\Longleftrightarrow\quad f_{1}(a)=f_{1}(a^{\prime}).$
They are functionally $k$-equivalent w.r.t. $D$ (denoted $a\equiv^{k}_{D}a^{\prime}$) when there exist $k$ distinct contexts $c_{1},\dots,c_{k}\in{\mathcal{X}}$ such that for every $r\leq k$ both $(a,c_{r})$ and $(a^{\prime},c_{r})$ appear in $D$ and $f(a,c_{r})=f(a^{\prime},c_{r})$. We call each $(a,c_{r}),(a^{\prime},c_{r})$ an evidence pair.
The coverage principle by itself is only a necessity statement: outside the $k$-coverage region a purely pattern matching learner’s predictions are unconstrained. To convert this into a sufficient data condition we adopt an explicit inductive bias, matching the premise of Result 5.1:
Learner assumption: Whenever two fragments become linked by $k$ independent (i.e., pairwise from distinct contexts) evidence pairs, the learner treats them as functionally equivalent; the learner also propagates equivalence transitively along chains of such links.
With this rule in place the relevant structure inside each class $E_{b}$ is the $k$-evidence graph: vertices are the $|{\mathcal{X}}|$ first-hop pairs in the class and an edge connects two vertices whenever the pair is observed at least $k$ times in shared contexts. If that graph is connected (every vertex reachable from every other), then every fragment in $E_{b}$ is linked by a chain of $k$-evidence steps and is therefore recognized as equivalent by the learner. Hence
Data sufficiency criterion: If the $k$-evidence graph of each class $E_{b}$ is connected, the learner generalizes perfectly to all in-domain (ID) inputs.
This criterion requires connectivity, and we aim to derive the condition to yield the connectivity with high probability using Erdős–Rényi model [96, 97]. The minimal dataset size achieving this with high probability is denoted $N_{\mathrm{req}}(|{\mathcal{X}}|,k)$.
Note that the $k$-evidence graph on $E_{b}$ is the restriction of the substitution graph $\mathcal{G}_{D,k}$ (Def. 3.2) to the vertex set $E_{b}\times\{x_{3}\}$ for any fixed $x_{3}$. Connectivity of every class therefore implies that every first-hop fragment lies in the same connected component as some training input, i.e., the entire in-domain set is contained in $k$-coverage. Under the learner assumption, this is both necessary and sufficient for perfect ID generalization, yielding Res. 5.1.
E.1 Step 1: probability of a single evidence pair
Evidence pair probability
Fix two distinct first–hop fragments $i,j\in E_{b}$ and a context $c\in{\mathcal{X}}$. We want to find $p_{1}$, the probability that context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$:
$p_{1}:=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
Let $q:=1/|{\mathcal{X}}|^{3}$ denote the probability of drawing any specific triple in a single draw. Using the inclusion–exclusion principle:
$p_{1}$ $=\Pr[(i,c)\in D\text{ and }(j,c)\in D]$
$=1-\Pr[(i,c)\not\in D]-\Pr[(j,c)\not\in D]+\Pr[(i,c)\not\in D\text{ and }(j,c)\not\in D]$
$=1-(1-q)^{N}-(1-q)^{N}+(1-2q)^{N}$ (S1.1)
For $q\ll 1$ (which holds when $|{\mathcal{X}}|$ is large), we can use Taylor expansion:
$(1-q)^{N}$ $=1-Nq+\frac{N(N-1)}{2}q^{2}+O(q^{3})$
$(1-2q)^{N}$ $=1-2Nq+\frac{N(N-1)}{2}(2q)^{2}+O(q^{3})$ (S1.2)
Substituting these approximations:
$p_{1}$ $=1-2\left(1-Nq+\frac{N(N-1)}{2}q^{2}\right)+\left(1-2Nq+2N(N-1)q^{2}\right)+O(q^{3})$
$=1-2+2Nq-N(N-1)q^{2}+1-2Nq+2N(N-1)q^{2}+O(q^{3})$
$=N(N-1)q^{2}+O(q^{3})$
$=\frac{N(N-1)}{|{\mathcal{X}}|^{6}}+O\left(\frac{1}{|{\mathcal{X}}|^{9}}\right)$
$=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))$ (S1.3)
Therefore, in the regime of interest ($N\gg|{\mathcal{X}}|$ but $N\ll|{\mathcal{X}}|^{3}$):
$\boxed{p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}(1+o(1))}$
Equation (E.1) gives an exact expression for the probability (up to lower-order terms) that the single context $c$ provides an evidence pair for the functional equivalence of fragments $i$ and $j$.
Remark on “lucky coincidences”
Because functional $k$-equivalence demands consistency across all $k$ evidences, a single coincidental equality $f(i,c)=f(j,c)$ with $i\not\sim j$ can only masquerade as evidence when $k=1$ and the dataset lacks any contradicting context $c^{\prime}$. For $k\geq 2$ the joint probability that two independent contexts simultaneously produce such coincidences is $|{\mathcal{X}}|^{-k}$ per fragment pair and hence negligible relative to the isolate probability once $N=\Omega(|{\mathcal{X}}|^{2})$. Consequently, the effects of such “lucky coincidences” on the lower bound in Res. 5.1 can be neglected.
E.2 Step 2: probability of observing $k$ evidences for one fixed pair
Having established the probability $p_{1}$ for a single evidence pair in Step 1, we now derive $p_{k}$, the probability that a dataset provides at least $k$ distinct contexts as evidence for the functional equivalence of two fragments $i$ and $j$. This probability will determine the edge probability in the random graph model analyzed in Step 3.
Indicators for one fragment pair
Fix two distinct first-hop fragments $i,j\in E_{b}$ and, for each context (third token) $c\in{\mathcal{X}}$, set
@@ E.2 Step 2: probability of observing $k$ evidences for one fixed pair
Thus $Z_{ij}(c)=1$ exactly when the single context $c$ supplies an evidence pair for the functional equivalence of $i$ and $j$.
Single-context success probability
From (E.1),
$p_{1}:=\Pr[Z_{ij}(c)=1]=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)},\qquad(|{\mathcal{X}}|\to\infty).$
Negatively correlated counts and an i.i.d. surrogate
Because all $N$ draws come from a single multinomial $\bigl{(}N,1/|{\mathcal{X}}|^{3}\bigr{)}$, the indicators $\{Z_{ij}(c)\}_{c\in{\mathcal{X}}}$ are negatively correlated: drawing many triples with one context leaves fewer draws for the others. Negative correlation decreases the probability that several $Z_{ij}(c)$ equal 1 simultaneously. Hence the tail probability for the true count
$Y_{ij}:=\sum_{c\in{\mathcal{X}}}Z_{ij}(c)$
is upper-bounded by the tail of an i.i.d. binomial variable with the same single-trial success probability $p_{1}$. Concretely, define
$Y_{ij}^{\star}\sim\mathrm{Binom}\bigl{(}|{\mathcal{X}}|,p_{1}\bigr{)},\qquad p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{6}}\bigl{(}1+o(1)\bigr{)}.$
Then for every real $t$
$\Pr\bigl{[}Y_{ij}\geq t\bigr{]}\leq\Pr\bigl{[}Y_{ij}^{\star}\geq t\bigr{]}.$
Using $Y_{ij}^{\star}$ therefore overestimates the chance of obtaining $k$ or more distinct evidence contexts, which is conservative for our goal of deriving a lower bound on the required dataset size $N$.
Since $Y_{ij}^{\star}$ is binomial with mean
$\mu:=\mathbb{E}[Y_{ij}^{\star}]=|{\mathcal{X}}|\,p_{1}=\frac{N^{2}}{|{\mathcal{X}}|^{5}}\bigl{(}1+o(1)\bigr{)},$
we may work henceforth with $Y_{ij}^{\star}$ alone; the resulting bounds apply verbatim to the original $Y_{ij}$.
Poisson tail via Le Cam

### v2

@@ D.3 Token Set Size Ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig.˜9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig.˜10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
@@ D.4 Task Ablation
Appendix E Partial Computation Observation Drives the Alignment of Functional Equivalence Representation and Vocabulary Space
In this section, we investigate how exposure to partial computations affects the interpretability of intermediate state representations through vocabulary space alignment. We compare two training conditions on a modified 2-Hop task with $\left|{\mathcal{X}}\right|=50$ and $N=10\text{k}$, after 40k epochs of training:
  1. 1.
  2. 2.
To assess interpretability, we measure the Mean Reciprocal Rank (MRR) of intermediate state representations when projected to vocabulary space using the unembedding matrix. The low MRR indicates that the model’s internal representation of intermediate state $b$ aligns with the corresponding vocabulary token.
Fig.˜15 shows a striking contrast between the two conditions. Under standard training, the MRR score remains very high throughout training, indicating that intermediate representations are not aligned with vocabulary space despite the model successfully learning the compositional task. However, when partial computations are included, the MRR score becomes very high, demonstrating clear vocabulary alignment.
This finding has important implications for mechanistic interpretability research: the absence of interpretable representations through the logit lens does not indicate the absence of structured internal computation. Furthermore, it suggests that interpretability techniques may need to account for how training data shapes the alignment between internal representations and vocabulary space, rather than assuming such alignment emerges naturally from task performance.
Appendix F Proofs: Asymptotically Perfect Coverage in 2-Hop Task
We establish a scaling-law analysis for the 2-Hop task (as illustrated in Fig.˜2(a)) within our formal framework for understanding pattern matching (Sec.˜3). While a simplified statement (Theorem˜6.1) is already presented in the main text, we provide the complete results of our analyses here in great detail, along with their rigorous and self-contained proofs. This appendix is long; we have organized it into the following structure.
  • In Sec.˜F.1, we first review and (re)define the problem setting for 2-Hop task. Since the functional ($k$-)equivalence at indices 1 and 2 matters for the 2-Hop task, we argue that it is enough for us to focus on $I=\{1,2\}$, thereby simplifying some key definitions (e.g., substitution graph, $k$-coverage) accordingly.
  • In Sec.˜F.2, we state and prove our main theorems that characterize the number of training data sufficient for a learner to achieve a perfect in-domain generalization with high probability. There are two main theorems: one is for $k\geq 2$ (Theorem˜F.7) and another is for $k=1$ (Theorem˜F.8). As a consequence, we further assume a specific regime of set cardinalities—namely, where all token sets have $\Theta(n)$ elements—and provide Corollary˜F.9, combining both cases for $k\geq 2$ and $k=1$.
    • We also claim that our results are nearly tight, especially for $k\geq 2$. It is done by constructing a worst-case subclass of 2-Hop problems in which, given a dataset slightly smaller than our upper bound, the learner fails in achieving a perfect ID generalization with high probability (Theorems˜F.16 and F.17).
    Theorem˜6.1 in the main text is indeed a simplified combination of Corollaries˜F.9 and F.17.
  • In Sec.˜F.3, we provide (almost) all postponed proofs of lemmas used in Sec.˜F.2. We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), which are the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
    • Sec.˜F.3.1 is dedicated to explaining the renowned Poissonization technique, rigorously proving one of our main lemmas, Lemma˜F.12. Under a usual fixed-sized dataset sampling scheme, the edge connections of the substitution graph are not guaranteed to be independent. Nevertheless, the Poissonization technique enables us to disentangle these dependencies, thereby simplifying our analysis.
    • Sec.˜F.3.2 reviews the binomial random $k$-intersection graphs, a topic in the random graph theory literature. We bring some results on the threshold functions for their connectivity (without bringing their involved proofs). It is crucial because we show that the $b$-evidence graphs can be regarded as binomial random $k$-intersection graphs under the Poissonization.
F.1 Preliminaries: Definitions, Assumptions, and Facts
Ground-truth mapping.
Consider a ground-truth mapping $f:{\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}\to Y$ defined as $f(x_{1},x_{2},x_{3})=f_{2}\left(f_{1}(x_{1},x_{2}),x_{3}\right)$, a two-hop composition of primitive functions $f_{1}:{\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\to B$ and $f_{2}:B\times{\mathcal{X}}_{3}\to Y$. Here, the set $B:=f_{1}({\mathcal{X}}_{1}\times{\mathcal{X}}_{2})$ is a collection of intermediate states, implicitly assuming surjectivity of $f_{1}$ without loss of generality.
Train dataset.
We take a train dataset $D$ by independently sampling input sequences $N$ times from the uniform distribution over ${\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}$. For the sake of simplicity, we allow duplicates/replacements of train samples in $D$, thereby considering $D$ as a multiset.
We find it helpful in our analysis to define a random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ that has a one-to-one correspondence with a dataset constructed by sampling with replacement as described below.
Proposition F.1.
Given a train dataset $D$ and for each ${\bm{x}}\in{\mathbb{X}}$, define $Z_{\bm{x}}$ as the sampled count of $\bm{x}$} (i.e., the number of identical duplicates equal to $\bm{x}$}) in $D$. Then, if $D$ is generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$). That is, for all collections of integers $z_{\bm{x}}\geq 0$ ($\forall{\bm{x}}\in{\mathbb{X}}$) such that $\sum_{{\bm{x}}\in{\mathbb{X}}}z_{\bm{x}}=N$,
${\mathbb{P}}(Z_{\bm{x}}=z_{\bm{x}}\;\forall{\bm{x}}\in{\mathbb{X}})=\frac{N!}{|{\mathbb{X}}|^{N}\cdot\prod_{{\bm{x}}\in{\mathbb{X}}}(z_{\bm{x}}!)}.$
The one-to-one correspondence between $D$ and $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ is straightforward: their realizations uniquely determine each other (up to the order of sampling).
A learner making predictions using functional $k$-equivalence.
Let us consider an artificial learner that is perfectly trained on $D$ (i.e., always predicts correctly on $D$). We suppose that the learner confidently predicts the output for the elements in $k$-coverage of $D$ (defined with functional $k$-equivalence), while producing a random (or unconfident) prediction outside of it. However, the confident prediction is not necessarily correct in general. Luckily, since we suppose a clear two-hop compositional structure of the task, the confident predictions based on functional equivalences at indices $I=\{1,2\}$ are guaranteed to be correct. Hence, to study the scope of pattern-matching-based in-domain (ID) generalization as confident and correct predictions on unseen inputs, we restrict the scope of our analysis to the index set $\{1,2\}$ for our analysis in this appendix.
To formalize how far the learner can generalize (Def.˜F.4) and how far we hope the learner to generalize (Def.˜F.2), let us (re)define the key terms for our analysis of the two-hop task, based on the restriction with respect to the index set $I=\{1,2\}$. For brevity, let us write $x_{12}:=(x_{1},x_{2})\in{\mathcal{X}}_{12}:={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}$.
Definition F.2 (In-domain closure, in terms of $I=\{1,2\}$).
For a train dataset $D$, its in-domain closure $\overline{D}$ is defined as
$\overline{D}=\left\{(x_{12},x_{3})\!\in\!{\mathbb{X}}:\exists(x^{\prime}_{12},x^{\prime}_{3})\!\in\!{\mathbb{X}}\text{ such that }f_{1}(x_{12})\!=\!f_{1}(x^{\prime}_{12}),(x^{\prime}_{12},x_{3})\!\in\!D,\text{and }(x_{12},x^{\prime}_{3})\!\in\!D\right\}.$
Also, its element $(x_{12},x_{3})\in\overline{D}$ is said to be an in-domain data.
According to the definition above, the in-domain closure $\overline{D}$ contains all input sequences $(x_{12},x_{3})$ such that (i) its subsequences $x_{12}$(${\bm{x}}_{I}$) and $x_{3}$(${\bm{x}}_{I^{\mathsf{c}}}$) has already been observed in the train dataset $D$, and (ii) its output can be inferred using functional equivalence at indices $I=\{1,2\}$, namely, $f(x_{12},x_{3})=f(x^{\prime}_{12},x_{3})$ for a training data $(x^{\prime}_{12},x_{3})\in D$. Hence, the set $\overline{D}$ is the largest set of input sequences whose output can be correctly predicted using the exact task structure (e.g., the functional equivalence at $\{1,2\}$). We also remark that it is obvious to see that $D\subset\overline{D}$. We wish the learner would acquire such prediction capabilities from a training dataset. Nevertheless, the task structure is not directly accessible to our learner since it can only identify the necessary functional relationships through functional $k$-equivalences (Def.˜3.1).
On the other hand, the $k$-coverage quantifies how far the assumed learner can actually make correct predictions using the observed functional $k$-equivalences from the training dataset $D$. The definition (Def.˜F.4) resembles that of the connectivity of a graph with a vertex set ${\mathcal{X}}_{12}$.
Recall that in our main text, we defined a substitution graph with the entire input space $\mathbb{X}$} as its vertex set since we aim to check all functional equivalences across all index sets. However, since we are particularly interested in the functional equivalences at $I=\{1,2\}$, we can simplify its definition as below.
Definition F.3 (Substitution graph, in terms of $I=\{1,2\}$).
For a train dataset $D$ and a positive integer $k$, we define a substitution graph (in terms of $I=\{1,2\}$) as ${\mathcal{G}}_{\bullet}^{(D,k)}=(V_{\bullet},E_{\bullet}^{(D,k)})$. Here, the vertex set is $V_{\bullet}:={\mathcal{X}}_{12}$ and the edge set is
$E_{\bullet}^{(D,k)}:=\left\{\left\{x_{12},x^{\prime}_{12}\right\}\subset V_{\bullet}:x_{12}\neq x^{\prime}_{12}\text{ and }\left|S_{f}(x_{12},x^{\prime}_{12}\mid D)\right|\geq k\right\},$
where the set $S_{f}(x_{12},x^{\prime}_{12}\mid D)$ is defined as
$S_{f}(x_{12},x^{\prime}_{12}\mid D):=\left\{\bar{x}_{3}\in{\mathcal{X}}_{3}:f(x_{12},\bar{x}_{3})=f(x^{\prime}_{12},\bar{x}_{3}),(x_{12},\bar{x}_{3})\in D,(x^{\prime}_{12},\bar{x}_{3})\in D\right\}.$
In other words, in the substitution graph ${\mathcal{G}}_{\bullet}^{(D,k)}$, two vertices $x_{12}$ and $x^{\prime}_{12}$ are adjacent if they have at least $k$ distinct co-occerences in $D$. With this definition, we define the $k$-coverage as follows.
Definition F.4 ($k$-coverage, in terms of $I=\{1,2\}$).
For a train dataset $D$ and a positive integer $k$, the $k$-coverage of $D$ (in terms of $I=\{1,2\}$) is a subset of $\mathbb{X}$} defined as
$\mathrm{Cover}_{k}(D)$
@@ F.1 Preliminaries: Definitions, Assumptions, and Facts
$\!=\!\left\{\!(u_{0},x_{3})\!\in\!{\mathbb{X}}:\exists\ell\!\geq\!0,\,\exists u_{0},u_{1},\cdots,u_{\ell}\!\in\!{\mathcal{X}}_{\!12}\text{ such that}\left\{u_{i\!-\!1},u_{i}\right\}\!\in\!E_{\bullet}^{(D,k)}\text{ ($\forall i\!\in\![\ell]$) \& }(u_{\ell},\!x_{3})\!\in\!D\!\right\}\!,$
where $\ell]=\{1,2,3,\cdots,\ell\}$. Also, an element $(x_{12},x_{3})\in\mathrm{Cover}_{k}(D)$ is said to be covered.
Again, we can easily verify that $D\subset\mathrm{Cover}_{k}(D)$ by choosing $\ell=0$ in the definition above. However, the inclusion relation between the in-domain coverage ($\overline{D}$) and the $k$-coverage ($\mathrm{Cover}_{k}(D)$) cannot be explicitly determined in general. From the next sub-section, we aim to analyze when (or, with how large $D$) we do have a relation $\overline{D}\subset\mathrm{Cover}_{k}(D)$”, which we call perfect coverage of in-domain data, with high probability (with respect to the randomness of $D$).
We additionally introduce some useful symbols for our later proofs. We denote by $B_{D}$ the set of all intermediate states observed in $D$: i.e.,
$B_{D}:=\left\{b\in B:\exists(x_{12},x_{3})\in D\text{ such that }f_{1}(x_{12})=b\right\}.$
Moreover, we introduce the $b$-evidence graph ${\mathcal{G}}_{b}^{(D,k)}$, defined as a subgraph of ${\mathcal{G}}_{\bullet}^{(D,k)}$ induced by the set $V_{b}=f_{1}^{-1}(\{b\}):=\left\{x_{12}\in{\mathcal{X}}_{12}:f_{1}(x_{12})=b\right\}\subset V_{\bullet}$ of vertices sharing the same intermediate state $b\in B$.
Definition F.5 (Evidence graphs).
For a train dataset $D$, a positive integer $k$, and an intermediate state $b\in B$, we define a $b$-evidence graph (in terms of $I=\{1,2\}$) as ${\mathcal{G}}_{b}^{(D,k)}=(V_{b},E_{b}^{(D,k)})$. Here, the vertex set is $V_{b}:=f_{1}^{-1}(\{b\})$ and the edge set is
$E_{b}^{(D,k)}:=\left\{\left\{x_{12},x^{\prime}_{12}\right\}\subset V_{b}:x_{12}\neq x^{\prime}_{12}\text{ and }\left|S(x_{12},x^{\prime}_{12}\mid D)\right|\geq k\right\},$ (5)
where the set $S(x_{12},x^{\prime}_{12}\mid D)$ is defined as
$S(x_{12},x^{\prime}_{12}\mid D):=\left\{\bar{x}_{3}\in{\mathcal{X}}_{3}:(x_{12},\bar{x}_{3})\in D,(x^{\prime}_{12},\bar{x}_{3})\in D\right\}.$ (6)
Lastly, we introduce an assumption that will be assumed only when we claim the tightness of our sample complexity upper bound results.
Assumption F.6.
The learner uses the parameter $k\geq k_{\star}+1$, where $k_{\star}\geq 0$ is defined as
$k_{\star}=\max_{(b,b^{\prime})\in B^{2}}\left|H(b,b^{\prime})\right|\qquad\text{subject to}\quad b\neq b^{\prime},$
and $H(b,b^{\prime})=\left\{x_{3}\in{\mathcal{X}}_{3}:f_{2}(b,x_{3})=f_{2}(b^{\prime},x_{3})\right\}$.
According to the assumption above, $f_{2}(b,x_{3})=f_{2}(b^{\prime},x_{3})$ can hold for more than $k_{\star}$ elements in ${\mathcal{X}}_{3}$ only when $b=b^{\prime}$. It is quite a natural assumption for general 2-Hop tasks since it naturally eliminates pathological examples that can barely be regarded as 2-Hop tasks (e.g., a constant function $f$ on $\mathbb{X}$}).
The assumption above additionally assumes that our learner pessimistically recognizes the functional equivalences between $\{1,2\}$-subsequences, by using $k\geq k_{\star}+1$. If $k=k_{\star}+1$, the learner is using the precise value of $k$ to recognize functionally equivalent subsequences sharing the same intermediate state; a higher value of $k$ indicates a more pessimistic learner. Because of this pessimism, once the learner recognizes that $x_{12}$ and $x^{\prime}_{12}$ are functionally $k$-equivalent (i.e., $\left\{x_{12},x^{\prime}_{12}\right\}\in E^{(D,k)}_{\bullet}$), it is guaranteed that $x_{12}$ and $x^{\prime}_{12}$ share the same intermediate state (i.e., $f_{1}(x_{12})=f_{1}(x^{\prime}_{12})$). More importantly, it implies that there must be no edges (in $E^{(D,k)}_{\bullet}$) connecting two graphs ${\mathcal{G}}^{(D,k)}_{b}$ and ${\mathcal{G}}^{(D,k)}_{b^{\prime}}$ for distinct $b,b^{\prime}\in B$; otherwise, there must be a pair of subsequences $x_{12}\in V_{b}$ and $x^{\prime}_{12}\in V_{b^{\prime}}$ that are functionally $k$-equivalent, contradicting the assumption.
Even though Assumption˜F.6 appears to pertain to the learning rule of the learner, it can also be interpreted as an assumption about the ground-truth mapping, which necessitates $k_{\star}\leq k-1$ for a given $k\geq 1$.
F.2 Main Theorem: Sample Complexity Bound for Perfect Coverage With High Probability
The ultimate goal of this appendix is to characterize a sufficient number of training data $|D|$ (i.e., sample complexity upper bound) so that the $k$-coverage includes all in-domain data, i.e., $\overline{D}\subset\mathrm{Cover}_{k}(D)$, with high probability. Here, we will state and prove our main theorems, Theorem˜F.7 (for $k\geq 2$) and Theorem˜F.8 (for $k=1$), which are followed by a simplified combination of them (Corollary˜F.9).
Soon after, we will also claim the near-tightness of these theorems by showing that, for $k\geq 2$, the training datasets slightly smaller than the obtained sample complexity are sufficient to guarantee $\overline{D}\not\subset\mathrm{Cover}_{k}(D)$ with high probability (Theorems˜F.16 and F.17). Although we present the proofs of the theorems here, the complete proofs of most lemmas used in the main proofs are postponed to the later part of this appendix (Sec.˜F.3).
Here are the statements of our main theorems for the sample complexity upper bounds, which will be proved soon after.
Theorem F.7 (Sample Complexity Upper Bound, $k\geq 2$, $\downarrow$).
Under the two-hop task setup described in Sec.˜F.1, suppose that $k\geq 2$ is a universal constant. Let $\hat{n}:=\min_{b\in B_{D}}\left|V_{b}\right|$ and $\check{n}:=\max_{b\in B_{D}}\left|V_{b}\right|$, and assume that $\left|{\mathcal{X}}_{3}\right|=\Omega(\check{n})$. Fix any $\delta>0$. Then, there exists $N_{\delta}>0$ satisfying $N_{\delta}\leq O\left(\max\left\{\left(\frac{k!\,(\ln\hat{n})}{\hat{n}}\right)^{\!\!\frac{1}{2k}}\!\cdot\left|{\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\right|\!\sqrt{\left|{\mathcal{X}}_{3}\right|},\;\ln\left(\frac{1}{\delta}\right)\right\}\right)$ such that, for any large enough $\hat{n}$ and any integer $N>N_{\delta}$, we have ${\mathbb{P}}_{N}\left(\overline{D}\subset\mathrm{Cover}_{k}(D)\right)\geq 1-\delta.$ Here, ${\mathbb{P}}_{N}$ is the probability measure for the train dataset $D$ constructed by uniformly randomly sampling its elements NN times, allowing duplicates, independently from ${\mathbb{X}}={\mathcal{X}}_{1}\times{\mathcal{X}}_{2}\times{\mathcal{X}}_{3}$.
Theorem F.8 (Sample Complexity Upper Bound, $k=1$, $\downarrow$).